### v1

# Kibble-Zurek behavior in a topological phase transition with a quadratic band crossing

Abstract
Kibble-Zurek (KZ) mechanism describes the scaling behavior when driving a system across a continuous symmetry-breaking transition. Previous studies have shown that the KZ-like scaling behavior also lies in the topological transitions in the Qi-Wu-Zhang model (2D) and the Su-Schrieffer-Heeger model (1D), although symmetry breaking does not exist here. Both models with linear band crossings give that $\nu=1$ and $z=1$. We wonder whether different critical exponents can be acquired in topological transitions beyond linear band crossing. In this work, we look into the KZ behavior in a topological 2D checkerboard lattice with a quadratic band crossing. We investigate from dual perspectives: momentum distribution of the Berry curvature in clean systems for simplicity, and real-space analysis of domain-like local Chern marker configurations in disordered systems, which is a more intuitive analog to conventional KZ description. In equilibrium, we find the correlation length diverges with a power $\nu\simeq 1/2$. Then, by slowly quenching the system across the topological phase transition, we find that the freeze-out time $t_{\mathrm{f}}$ and the unfrozen length scale $\xi(t_{\mathrm{f}})$ both satisfy the KZ scaling, verifying $z\simeq 2$. We subsequently explore KZ behavior in topological phase transitions with other higher-order band crossing and find the relationship between the critical exponents and the order. Our results extend the understanding of the KZ mechanism and non-equilibrium topological phase transitions.
I Introduction
Quantum phase transitions are one of the fascinating topics in modern physics. Conventional continuous phase transitions are driven by spontaneous symmetry breaking and are usually described in terms of local order parameters that take continuous values Miransky (1994); Vojta (2003); Sachdev (2011); Nie and Zheng (2023a). In contrast, topological phase transitions are a new class of transitions beyond Landau’s symmetry breaking theory, characterized by a quantized, non-local topological invariant Hasan and Kane (2010); Qi and Zhang (2011). Recently, the non-equilibrium dynamics of topological phase transitions have received widespread attention both experimentally and theoretically Fläschner et al. (2018); D’Alessio and Rigol (2015); Hauke et al. (2014); Ye and Li (2020); Sun et al. (2018); Nie and Zheng (2023b), particularly concerning the critical behavior of quenching across topological phase transitions Liou and Yang (2018); Ulčakar et al. (2019, 2020); Sun et al. (2022); Kuo et al. (2021).
Relating conventional continuous phase transitions and nonequilibrium dynamics, the Kibble-Zurek (KZ) mechanism was originally proposed by Kibble in his study of the formation of topological defects in the early universeKibble (1976), and it has been applied by Zurek to study phase transitions in condensed matter systems Zurek (1985); Zurek et al. (2005). It elucidates the nonequilibrium properties of a system when it is quenched across a continuous phase transition in a finite time $\tau$, which shows universal scaling. In equilibrium, the theory of critical phenomena states that as the control parameters $V$ are tuned closer to their critical values $V_{\mathrm{c}}$, the correlation length $\xi$ and relaxation time $t_{\mathrm{r}}$ tend to diverge as $t_{\mathrm{r}}\sim|V-V_{\mathrm{c}}|^{-z\nu}$ and $\xi\sim|V-V_{\mathrm{c}}|^{-\nu}$, where $z$ is dynamical exponent and $\nu$ is spatial critical exponent. As the system slowly quenches across the transition, the relaxation time diverges as it approaches the critical point, and the system’s evolution ceases to be adiabatic. The moment when the system no longer evolves adiabatically is referred to as the freeze-out time $t_{\mathrm{f}}$. The freeze-out time and the unfrozen correlation length satisfy $t_{\mathrm{f}}\sim\tau^{z\nu/(1+z\nu)}$ and $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}$. The KZ behavior has been observed in diverse experimental systems, such as cosmic microwave background Bevis et al. (2008), superfluids Chesler et al. (2015); Mathey and Polkovnikov (2010), superconductor Sonner et al. (2015); Garaud and Babaev (2014); Monaco et al. (2006), liquid crystals Chuang et al. (1991), and ultracold atoms Donner et al. (2007); Weiler et al. (2008); Ko et al. (2019); Chen et al. (2011); Lamporesi et al. (2013); Navon et al. (2015); Clark et al. (2016); Keesling et al. (2019); Yi et al. (2020); Huang et al. (2021); Liang et al. (2022).
Although topological transitions can not be described by Landau’s symmetry breaking theory, the KZ-like behavior has already been introduced to topological systems. It is first proposed that effective correlation length can be extracted from Berry curvature profile in momentum space, and it shows power-law divergence at the transition Chen et al. (2017). However, this momentum-space description is not quite intuitive in contrast with the conventional KZ mechanism. Recently, significant breakthroughs have been made in the real-space analysis. It has been found that the local Chern marker (LCM) will show domain-like configurations when the system is disordered Ulčakar et al. (2020). As long as the disorder is weak, the disorder-averaged normalized autocovariance function of LCM is independent of the disorder strength. So the autocovariance function can be regarded as an intrinsic property of the original topological system. Then, the correlation length can be determined in this purely real-space method. During the quench, the domain-like configurations are the analogy to conventional topological defects and the correlation length shows KZ behavior. This method is first proposed in two-dimensional Qi-Wu-Zhang (QWZ) insulators Ulčakar et al. (2020) and then extended to one-dimensional Su-Schrieffer-Heeger (SSH) models Sun et al. (2022) by changing LCM to local winding number. However, since these two models have similar linear dispersion while gap closing, they have yielded the same critical exponents, $\nu=1,z=1$. Therefore, we naturally wonder whether different critical exponents can be obtained.
In this work, we systematically investigate the KZ behavior in the topological transition of a 2D checkerboard model with quadratic band crossing. For every scaling studied, two perspectives are adopted: momentum-space Berry curvature in the clean system, and the real-space local Chern marker in the disordered system. In the former, we mainly focus on the radial distribution of Berry curvature. In the latter, we are concerned with the domain-like configurations of LCM and extract real-space correlation lengths.
Figure 1: (a). Checkboard lattice model. Red (Blue) points denote the A (B) sublattices. NN hopping strength along the black arrows direction (opposite direction) is $J-\mathrm{i}V$ ($J+\mathrm{i}V$), which breaks the time-reversal symmetry as long as $V\neq 0$. Red (Blue) solid lines represent the NNN hopping strength $J_{1}^{\prime}(J_{2}^{\prime})$. (b). Topological phase diagram. Chern number $|C|=1$ acquires an opposite sign to $V$. At the transition, it is a time-reversal invariant semimetal. (c). The band structure of our model is in different regimes. The colors denote the polarization $\langle\sigma_{z}\rangle$. The gap closes with a quadratic dispersion at the transition $V=0$. We can see the band-inversion surface (BIS) (shown as the white curves) is not a circle around symmetric momentum as in the QWZ model Zhang et al. (2018); Sun et al. (2018), which is a fundamental effect induced by the quadratic band crossing.
We find that the length scale, in both perspectives, diverges near the critical point of the topological phase transition in equilibrium, satisfying $\xi\sim|V-V_{\mathrm{c}}|^{-1/2}$. By slowly quenching through the topological phase transition over a quench time $\tau$, we observe that the freeze-out time, defined by the delay time of the relaxation of Hall response in the clean system or the divergence of correlation length in the disordered system, satisfies $t_{\mathrm{f}}\sim\tau^{1/2}$. The corresponding length scale at the freeze-out time $t_{\mathrm{f}}$ satisfies $\xi(t_{\mathrm{f}})\sim\tau^{1/4}$, where the dynamic exponent $z=2$ and spatial critical exponent $\nu=1/2$ differs from the results of linear band crossings. Additionally, we extend the analysis to higher-order dispersion relations beyond quadratic band crossings and find that the KZ power-law relations still hold. The corresponding critical exponents are related to the band-crossing relations.
II Model
Two-dimensional (2D) checkerboard lattice Sun et al. (2009, 2011) is an example of a topological model with quadratic band crossing. We consider a simplified checkerboard lattice model consisting of two sublattices (A and B) with real and isotropic nearest-neighbor (NN) hoppings ($J$), two kinds of next-nearest-neighbor (NNN) hoppings ($J_{1}^{\prime},J_{2}^{\prime}$), and complex NN hopping ($\mathrm{i}V$) leading to a topological nontrivial phase (see Fig.1(a)). The Hamiltonian can be written as:
$\begin{split}\hat{H}&=\sum_{\mathbf{r}}|x,y\rangle\langle x,y|\otimes\begin{pmatrix}0&J-\mathrm{i}V\\ J+\mathrm{i}V&0\end{pmatrix}\\ &+\sum_{\mathbf{r}}\left[|x,y\rangle\langle x+1,y|\otimes\begin{pmatrix}J_{2}^{\prime}&0\\ J-\mathrm{i}V&J_{1}^{\prime}\end{pmatrix}+\text{H.c.}\right]\\ &+\sum_{\mathbf{r}}\left[|x,y\rangle\langle x,y+1|\otimes\begin{pmatrix}J_{1}^{\prime}&J+\mathrm{i}V\\ 0&J_{2}^{\prime}\end{pmatrix}+\text{H.c.}\right]\\ &+\sum_{\mathbf{r}}\left[|x,y+1\rangle\langle x+1,y|\otimes\begin{pmatrix}0&0\\ J+\mathrm{i}V&0\end{pmatrix}+\text{H.c.}\right]\\ \end{split}$ (1)
where $\mathbf{r}=(x,y)$ denote the unit cell position (unit cell is shown as a dashed ellipse in Fig.1(a)), and the $2\times 2$ matrices represent the sublattice hopping corresponding to the unit cell hopping.
After Fourier transformation, we can express the Hamiltonian in momentum space ^$\hat{\mathcal{H}}_{\mathbf{k}}$ with momentum $\mathbf{k}=(k_{x},k_{y})$ as
^$\begin{split}\hat{\mathcal{H}}_{\mathbf{k}}=&\left[-\left(J_{1}^{\prime}+J_{2}^{\prime}\right)\left(\cos k_{x}+\cos k_{y}\right)\right]\hat{I}\\ &-4J\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\right)\hat{\sigma}_{x}-2V\left(\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\right)\hat{\sigma}_{y}\\ &-\left(J_{1}^{\prime}-J_{2}^{\prime}\right)\left(\cos k_{x}-\cos k_{y}\right)\hat{\sigma}_{z}\end{split}$ (2)
where $\hat{I}$ and $\hat{\sigma}_{x,y,z}$ are the identity and Pauli matrices, respectively. We set $J=1$, $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}$ and periodic boundary condition for entire text afterwards. It is easy to read from Eq.2 that the system is an insulator without time-reversal symmetry as long as $V\neq 0$. Actually, the system has two topologically non-trivial phases, whose Chern number $C$ is determined by the sign of complex hopping $V$, as shown in Fig.1(b). Additionally, the energy band structures in different regimes are depicted in Fig.1(c). We can see that the energy gap closes at $\mathbf{G}=(k_{x},k_{y})=(\pi,\pi)$ when $V=V_{\mathrm{c}}=0$, exhibiting a quadratic band crossing. This is totally different from the QWZ model. There, the gap closes at a Dirac cone, and every Dirac cone changes the Chern number with $1$. Besides, the band-inversion surface (defined as $\langle\sigma_{z}\rangle=0$ contour) in the checkerboard lattice model forms a crisscross, while in the QWZ model, it is a circle around the band crossing point.
III Ground state
For conventional second-order phase transitions driven by spontaneous symmetry breaking, the correlation lengths exhibit a universal scaling behavior as the systems approach the critical point. When it comes to topological transitions, such as the QWZ model and the SSH model, a scaling of $\nu=1$ has been found. However, can this exponent be different in other models with different band crossings?
In the following, we investigate the checkerboard model from two perspectives: the momentum space in the clean system, and the real space in the disordered system.
III.1 Clean system
Figure 2: (a). Correlation length in the ground state could be determined by the FWHM of the radial distribution of Berry curvature with $J^{\prime}=1$. The closer to the transition, the width of the peak decreases. (b). This correlation length diverges in a power $\nu\simeq 1/2$ as the system approaches the phase transition.
Some studies have provided correlation functions for topological phases, defined as the Fourier transform of the Berry curvature Chen et al. (2017); Chen and Schnyder (2019), which measures the itinerant-circulation correlation between Wannier states. Thonhauser et al. (2005); Marzari et al. (2012). More directly, an alternative definition of correlation length can also be extracted from the profile of the Berry curvature in momentum space. Similar procedures have been used in studying the critical behavior in the QWZ insulators and the SSH model Ulčakar et al. (2020); Sun et al. (2022).
The occupied-band Berry curvature $F(\mathbf{k})$ can be expressed via Berry connection $A_{k_{\alpha}}=\left\langle u_{\mathbf{k}}\left|\mathrm{i}\partial_{k_{\alpha}}\right|u_{\mathbf{k}}\right\rangle$ ($\alpha=x,y$) as
$F(\mathbf{k})=\partial_{k_{x}}A_{k_{y}}-\partial_{k_{y}}A_{k_{x}}$ (3)
where $\left|u_{\mathbf{k}}\right\rangle$ is the occupied Bloch wavefunction at momentum $\mathbf{k}$. The Chern number mentioned before can be calculated by $C=\frac{1}{2\pi}\int_{\rm{FBZ}}F(\mathbf{k})\rm{d}^{2}\mathbf{k}$, where ”FBZ” represents the first Brillouin zone.
Then, by integrating the Berry curvature angularly, $f(k)=\frac{1}{k}\int_{|\mathbf{k}^{\prime}-\mathbf{G}|=k}F(\mathbf{k}^{\prime})\text{d}k^{\prime}$, we obtain its radial distribution $f(k)$. Fig 2(a) illustrates fixed $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}=1.5$ with six different values of complex NN hoppings strength $V$. It is evident that the radial distribution of the Berry curvature exhibits a peak. Initially, this peak width increases with the increasing $V$, reaching its maximum at the topological phase transition point $V=0$, and subsequently increases. We can find that there may exist critical behavior near the topological phase transition point.
To describe this critical behavior quantitatively, we define the correlation length $\xi_{k}$ by the inverse of the full width at half maximum (FWHM) value of the radial distribution of Berry curvature $f(k)$. A similar approach has been utilized in the QWZ insulators and the SSH model Ulčakar et al. (2020); Sun et al. (2022). As shown in Fig.2(b), we observe that as the system approaches the critical point $V=0$, the correlation length diverges. We plot the relationship between the correlation length and $V-V_{\mathrm{c}}$ in log-log coordinates under three sets of parameters, namely $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}=0.5$, $1$, $1.5$. We find that the correlation length satisfies $\xi_{k}\sim|V-V_{\mathrm{c}}|^{-\nu}$, where $\nu$ is approximately $1/2$. This critical exponent value is distinct from the case of linear dispersion, such as in the QWZ insulators where $\nu\simeq 1$.
Figure 3: (a). The real-space distribution of local Chern marker in a disordered sample with size $N\times N=50\times 50$. The area of the fragment domain increases to the order of the whole system sizes at the transition. (b). Divergence of real-space correlation length $\xi_{r}$ at transition with different system sizes in a log-log scale. Each data is an average of $20$ disorder samples. For a certain system size $N$, the correlation length shows good power law in $64<\xi<N/$ 6 regime (emphasized by the markers). The red line represent the $\left|V-V_{\mathrm{c}}\right|^{-1/2}$ scaling for ref. (c). The fitted power $\nu$ approaches $1/2$ as the system size increases. Here, the NNN hopping strength is set to $J^{\prime}=1$ and the disorder strength to $\delta V=0.05$.
III.2 Disordered system
Then, we switch to another perspective: real-space analysis in disordered systems. Following the previous studies, correlation length can be characterized by the real-space distribution of the local Chern marker (LCM) when the system is weakly disordered, where the LCM $c(\mathbf{r})$ is calculated via Prodan et al. (2010); Bianco and Resta (2011); Ulčakar et al. (2020)
$c(\mathbf{r})=2\pi\mathrm{i}\mathrm{Tr}\left\{\hat{P}(\mathbf{r})\hat{P}_{\mathrm{val}}\left[-\mathrm{i}\left[\hat{x},\hat{P}_{\mathrm{val}}\right],-\mathrm{i}\left[\hat{y},\hat{P}_{\mathrm{val}}\right]\right]\right\}.$ (4)
Here $\hat{P}(\mathbf{r})$ is the local projector to the unit cell at $\mathbf{r}$, $\hat{P}_{\mathrm{val}}$ is the projector to the valence band (the lower half spectrum), and $\hat{x},\hat{y}$ are the position operators. The LCM has the properties that it is uniform and equal to the Chern number in clean systems. As long as the disorder is weak and takes the thermodynamic limit, the mean value of LCM is still the Chern number in its clean corresponding, and the local deviation from the mean value shows a domain-like configuration, whose amplitude is proportional to the disorder strength, but its shape is independentUlčakar et al. (2020). Besides, if there are enough disorder samples, the disorder-averaged normalized autocovariance function converges to a stable limit. Thus, the disorder-averaged normalized autocovariance function calculated from a weakly disordered system can be regarded as an intrinsic character of its clean corresponding.
We consider the disorder is added on the time-reversal breaking hopping term, $V\to V_{\mathrm{dis}}(r)$, for $V$ is the only relevant parameter that controls the transition, with an independent uniform distribution $V_{\mathrm{dis}}(r)\in[V-\delta V,V+\delta V]$ with $\delta V=0.05$. The real-space distribution of the LCM in a certain disordered sample with periodic boundary conditions is illustrated in Fig.3(a) for several values of $V$, ranging from topological regions with $C=1$, across the critical point to another topological phase with $C=-1$. We can see that the sizes of domains reach their maximum value (about half the whole system size) at the transition, which indicates the divergence of correlation length in the thermodynamic limit.
We identify the correlation length $\xi_{r}$ in the real space to the distance where the disorder-averaged autocovariance function of the LCM distribution drops to zero App . Fig.3(b) shows that $\xi_{r}$ exhibits power-law scaling as $V$ approaches the critical point in the intermediate correlation length regime, i.e. $64<\xi_{r}<N/$ 6. That is because we are dealing with a finite-size lattice model, the maximal $\xi_{r}$ can not exceed the system size (infrared limit), and the minimal $\xi_{r}$ must saturate to the order of one lattice constant (ultraviolet limit). Only the intermediate regime suffers less from these two limits.
As the system size increases, the estimated scaling exponent approaches $1/2$, which is consistent with the results obtained from the Berry curvature in the clean system. From the above discussion, we find that in a 2D checkerboard model with quadratic band crossing, the critical phenomena of the topological phase transition are present in both momentum space and real space, with a power-law exponent equal to half of the results of linear band crossings.
IV Quench dynamics
In this section, we investigate the dynamical behavior of topological phase transitions in the 2D checkerboard lattice. The system is quenched across the topological phase transition by linearly changing the complex NN hoppings strength $V$ with time $t$, expressed as $V(t)=-1/2+t/\tau$, where $\tau$ denotes the quench time, which corresponds to the change in the Chen number from $1$ to $-1$ as shown by the arrow in the Fig.1(b).
We still work in two perspectives and focus on the freeze-out time $t_{\mathrm{f}}$ and unfrozen correlation length $\xi(t_{\mathrm{f}})$.
IV.1 Clean system
IV.1.1 Freeze-out time $t_{\mathrm{f}}$
Figure 4: (a). The dynamics of non-equilibrium Hall response during the quench. These curves collapse in a rescaled time coordinate. We identify the freeze-out time $t_{\mathrm{f}}$ as the time when the Hall response crosses zero, $\tilde{\Sigma}_{xy}(\hat{t}=t_{\mathrm{f}})=0$. The inset shows curves in a non-rescale $\hat{t}$ coordinate. (b). Freeze-out time with different quench time in a log-log plot, $t_{\mathrm{f}}$ scales with approximately $\sim\tau^{1/2}$. Here, the NNN hopping strength is set to $J^{\prime}=1$.
To explore the temporal scaling in the dynamical process, we choose to study the Hall response during the non-equilibrium process Kuo et al. (2021). The non-equilibrium Hall response is given by the integral of the Berry curvature $F_{\mathrm{b}}(\mathbf{k},t)$ weighted by the occupation probability $P_{\mathrm{b},\mathbf{k}}(t)$ in the upper (lower) band Caio et al. (2016); Hu et al. (2016):
$\begin{split}\tilde{\Sigma}_{xy}(t)=\frac{1}{2\pi}\sum_{\mathrm{b}}\int_{\mathrm{FBZ}}\text{d}^{2}\mathbf{k}P_{\mathrm{b},\mathbf{k}}(t)F_{\mathrm{b}}(\mathbf{k},t)\end{split}$ (5)
Here $\mathrm{b}\in\left\{\mathrm{val},\mathrm{con}\right\}$ refer to the valence and conduction band. We computed the evolution of the Hall response during the quench process, as depicted in Fig.4(a). The Hall response deviates from its original quantized value and relaxes to a new quantized value after the quench. By transforming the coordinates with $\hat{t}=t-t_{\mathrm{c}}$, we set the phase transition time $t_{\mathrm{c}}=\tau/2$ as the zero point and the freeze-out time $t_{\mathrm{f}}$ is defined as the time at which the Hall response relaxes to the average of its initial and final values. We plot the relationship between the freeze-out time $t_{\mathrm{f}}$ and the quench time $\tau$ in the log-log scale and found that it follows a power-law relationship, with the corresponding exponent close to $1/2$. This matches the prediction of the KZ mechanism for linear quenching with $z\nu=1$, $t_{\mathrm{f}}\sim\tau^{\frac{z\nu}{1+z\nu}}\simeq\tau^{1/2}$.
IV.1.2 Unfrozen correlation length $\xi_{k}(t_{\mathrm{f}})$
We subsequently investigate the spatial scalings. The KZ mechanism predicts that at the freeze-out time, the correlation length follows a power-law relationship. By obtaining the evolved wave functions during the quenching process, we calculated the corresponding Berry curvature and extracted the correlation length. Fig.5(a) illustrates the radial distribution of the Berry curvature at corresponding freeze-out time $f(k,{t_{\mathrm{f}}})$ under different quench time. We observed that the width of the peak decreases with increasing quench time. Furthermore, the relationship between the extracted correlation length and the quench time is plotted for $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}=0.5,1,1.5$ and $J=1$. We found power-law relationships between the correlation length and quench time for different correlation lengths, with corresponding critical exponents very close to $1/4$. From the critical exponent of the correlation length at equilibrium, we obtain $\nu\simeq 1/2$. The power-law behavior of the freeze-out time gives the dynamic exponent $z\simeq 2$. According to the KZ mechanism, the unfrozen correlation length scales as $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}\simeq\tau^{1/4}$. Therefore, our results are consistent with the KZ prediction. This critical exponent with a quadratic band crossing is half of the linear band crossing Ulčakar et al. (2020); Sun et al. (2022).
Figure 5: (a). Radial distribution of the Berry curvature at the freeze-out time $t_{\mathrm{f}}$ with $J^{\prime}=1$. (b). Correlation length $\xi_{k}$ at the freeze-out time obtained by the Berry curvature has a scaling with quench time $\tau$, $\xi\sim\tau^{1/4}$.
IV.2 Disordered system
IV.2.1 Freeze-out time $t_{\mathrm{f}}$
It is not straightforward to generalize the Hall response to the disordered systems. We need real-space analysis to show analogs to the formation of topological defects. Therefore, we switch back to explore the LCM distribution.
The distribution of the LCM in quench dynamics can be calculated by replacing the projector onto the valence band $\hat{P}_{\mathrm{val}}$ with the projector onto the instant occupied subspace. The evolution of the LCM profile during the quench process with $\tau=10$ and $\tau=100$ under weak disorder is depicted in Fig.6(a). At the beginning of the slow quench, their LCM profiles are similar since they both follow the adiabatic evolution. For longer quench time, the system enters into the freeze-out regime later and holds a larger domain size after quench.
Figure 6: (a). The spatial distribution of the local Chern marker during a fast ($\tau=10$) and a slow ($\tau=100$) quench with size $N=50$. The slower the quench, the larger the post-quench domain size. (b). The dynamics of correlation length in real space in a rescaled coordinate. Curves collapse near the transition point. The freeze-out time $t_{\mathrm{f}}$ can be determined by the time delay of the trend of divergence in $\xi_{r}$ while crossing the transition. The inset shows the same data in a non-rescaled coordinate. Each data is an average from $10$ disorder samples. (c). Left panel: freeze-out time has a power-law scaling with the quench time, and the power-law exponent is close to $1/2$ for different system sizes. Right panel: estimated exponents with different system sizes. (d). Left panel: correlation length at the freeze-out time shows a scaling behavior, and the power-law exponent is close to $1/4$ for different system sizes. Right panel: estimated exponents with different system sizes. Here, the NNN hopping strength is set to $J^{\prime}=1$ and the disorder strength to $\delta V=0.05$.
Like before, we extract the evolution of the correlation length $\xi_{r}$ over different quenching time, as shown in Fig.6(b). During the evolution, the correlation length $\xi_{r}$ adiabatically follows the instant ground-state value initially. Near the transition, the system enters into the freeze-out regime. Although the correlation length still increases, the trend of divergence occurs later than the critical point, so it shows a maximum after the transition. Here, we claim this time delay could be identified as the freeze-out time $t_{\mathrm{f}}$ (We verify that this method is also applicable to clean systems as discussed in Appendix App ). After that, the correlation length is stabilized around a constant larger than the ground-state value.
In Fig.6(c), we observe that the freeze-out time $t_{\mathrm{f}}$ exhibits a power-law relationship with the quench time $\tau$, with approximately the same power-law exponent of $1/2$ as observed in the clean system.
IV.2.2 Unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$
Similarly, we extract the unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$ App from the LCM at the freeze-out time $t_{\mathrm{f}}$ to analyze. We plot the relationship between unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$ and the quench time on a log-log scale in the Fig.6(d). We find that $\xi_{r}(t_{\mathrm{f}})$ scales with $\tau$ in an exponent of approximately $1/4$. Correspondingly, the critical exponent of the correlation length at equilibrium yields $\nu\simeq 1/2$, and the power-law behavior of the freeze-out time gives the dynamic exponent $z\simeq 2$. According to KZ mechanism, the unfrozen correlation length is expected to scale as $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}\simeq\tau^{1/4}$. Thus, our findings are in agreement with the KZ prediction as observed in the clean system.
Therefore, in the checkerboard lattice with quadratic band crossing, when quenching across the topological phase transition, we observe temporal and spatial power-law behavior analogous to the Kibble-Zurek (KZ) mechanism in both momentum space and real space. The temporal power-law exponent is the same as that of linear band crossing, while the spatial power-law exponent is half of that.
V Outlook on higher-order band crossing
The critical exponent obtained from the quadratic band crossing differs from that of the previously studied linear band crossing. Therefore, it remains uncertain whether the KZ power-law scaling behaviors persist in systems with higher-order dispersion relations. If so, the corresponding critical exponent and its relationship with the band dispersion need further investigation. Accordingly, we can write down a general low-energy two-band effective model with higher-order crossing in momentum space, and explore the KZ behavior via methods mentioned in clean systems.
The low-energy effective Hamiltonian reads Chen et al. (2017)
$\begin{split}H_{p}(k,\phi)&=\eta_{p}k^{p}\cos(p\phi)\sigma_{x}+\eta_{p}k^{p}\sin(p\phi)\sigma_{y}+M\sigma_{z},\end{split}$ (6)
where $k$ and $\phi$ represent the distance and polar angle relative to band-crossing momentum, respectively and integer $p$ denotes its crossing order, with $p=1$ representing linear band crossing and $p=2$ representing quadratic band crossing, and so forth. The topological phase transition occurs at $M=0$ with the change in Chern number before and after the transition given by $C(M\textgreater 0)-C(M\textless 0)=p$. Based on this Hamiltonian, we can first compute the Berry curvature in the ground state and obtain the correlation length. We numerically verify that the correlation length in ground state satisfies $\xi\sim|M-M_{\mathrm{c}}|^{-\nu}$, where $\nu$ is approximately $1/p$.
Similarly to the previous analysis, we quench across the topological phase transition, measure the Hall response, and extract the power-law relationship between the freeze-out time and the quench time. We find that temporal scaling is independent of the order of band crossing with the corresponding exponent close to $1/2$. Similarly, by measuring the time evolution of the Berry curvature, we obtain the correlation length at the freeze-out time under different quench time, which also satisfies a power-law relationship. We observe that the corresponding exponents are distinct and close to $(2p)^{-1}$.
Figure 7: (a). Correlation length in the ground state diverges in a power $\nu\simeq 1/3$ for $p=3$ and $\nu\simeq 1/4$ for $p=4$ as the system approaches the phase transition. (b). Freeze-out time has a power-law scaling with the quench time, $t_{\mathrm{f}}\sim\tau^{1/2}$ with $p=3$ and $p=4$. (c). Correlation length at the freeze-out time shows a scaling behavior for $p=3$, $\xi(t_{\mathrm{f}})\sim\tau^{1/6}$ and $p=4$, $\xi(t_{\mathrm{f}})\sim\tau^{1/8}$.
VI Summary
In this work, we investigated the KZ behavior in a topological phase transition with a quadratic band crossing. We used dual perspectives: momentum space in clean systems and real space in disordered systems. Both perspectives give consistent scalings, which is distinct from previously studied linear band crossing models. Further, we go beyond quadratic to higher orders. They also follow KZ behavior but acquire different critical exponents.
Our study extends the KZ behavior in topological transitions beyond linear band crossing and confirms the consistency between correlation length from Berry curvature and LCM. For future studies, we get some hint from this work that the LCM correlation length may characterize the Anderson insulator transition if the disorder is strong.
Acknowledgements.
XN thanks Hui Zhai, Wei Zheng, Fan Yang and Hanteng Wang for their helpful discussions. This work is supported by NSFC (Grants No. GG2030007011 and No. GG2030040453) and Innovation Program for Quantum Science and Technology (Grants No. 2021ZD0302004). The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of University of Science and Technology of China. This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
Appendix A Extraction of correlation length from the local Chern marker
Figure 8: (a). The distribution of the LCM at freeze-out time for different quench time. (b). Disorder-averaged autocovariance function calculated from the LCM. The inset shows the same data in a non-rescaled coordinate. Here, the NNN hopping strength is set to $J^{\prime}=1$, and the average is performed over $10$ disorder realizations $\delta V=0.05$ and size $N=50$.
Fig.8(a) shows the real-space distribution of the local Chern marker (LCM) at freeze-out time for different quench time during the quench process. To estimate the size of inhomogeneities in the LCM, we calculated the normalized autocovariance function of the real-space distribution function $\delta c(r)$:
$R(r)=\frac{\sum_{|\mathbf{r}|=r}\sum_{\mathbf{r}^{\prime}}\delta c\left(\mathbf{r}^{\prime}\right)\delta c\left(\mathbf{r}+\mathbf{r}^{\prime}\right)}{\sum_{|\mathbf{r}|=r}\sum_{\mathbf{r}^{\prime}}\delta c\left(\mathbf{r}^{\prime}\right)^{2}},$ (7)
where $\delta c(\mathbf{r})=c(\mathbf{r})-\bar{c}$ is the LCM fluctuation. Similar to previous work, we determine the correlation length in the LCM as the distance at which the disorder-averaged autocovariance function crosses zero, $\bar{R}(\xi_{r})=0$. The autocovariance function of the LCM at freeze-out time for different quench time, is shown in Fig.8(b), plotted as a function of $r/\tau^{1/4}$. Since the correlation length at the freeze-out time satisfies $\xi_{r}\sim\tau^{1/4}$, The autocovariance functions of the LCM in Fig.8(b) overlap near $\bar{R}(\xi_{r})=0$.
Appendix B Alternative characterization of the freeze-out time in momentum space
Figure 9: (a). The dynamics of correlation length in momentum space in a rescaled coordinate. Curves collapse near the transition point. The freeze-out time $t_{\mathrm{f}}$ can be determined by the first peak after the transition. The inset shows the same data in a non-rescaled coordinate. (b). Freeze-out time $t_{\mathrm{f}}$ has a power-law scaling with the quench time $t_{\mathrm{f}}\sim\tau^{1/2}$. (c). Correlation length at freeze-out time $t_{\mathrm{f}}$ shows a scaling behavior $\xi(t_{\mathrm{f}})\sim\tau^{1/4}$.
In the main text, we use the nonequilibrium Hall response to determine the freeze-out time during the quench in clean systems. The freeze-out time by definition characterizes the time when the system’s relaxation rate matches the cooling rate. That is to say, it could be seen as a time delay when the system realizes the transition has occurred. So, the time delay of the trend of divergence of the correlation length can also be utilized to identify the freeze-out time $t_{\mathrm{f}}$. This method can be regarded as a supplement to the Hall response criterion and can be generalized to real space for disordered systems directly, while Hall response is not easy to find analog in real space.
In Fig.9, we extract the freeze-out time $t_{\mathrm{f}}$ from the first peak after the transition. It shows the scaling of the freeze-out time $t_{\mathrm{f}}$ to the quench time $\tau$. And the unfrozen correlation length $\xi_{k}(t_{\mathrm{f}})$ also follows KZ’s prediction.